OpenStudy (anonymous):

determine the pressure at the bottom of a column of mercury 70 cm high ( density of mercury = 13.6 g/cm^3 ) i want the final result in pascal please

OpenStudy (jamesj):

Use the formula I gave you on your last question.

OpenStudy (anonymous):

ok but its cm i want it in pascal

OpenStudy (jamesj):

So convert everything to SI units: distance in meters g in m/s^2 density is kg/m^3

OpenStudy (anonymous):

can you tell me the final result please i solve it

OpenStudy (jamesj):

Pressure = (Force)/(unit Area) = N / m^2 = kg.m/s^2 / m^2 = kg / (m.s^2) Now for h = height in meters rho = density in kg/m^3 g = acceleration in m/s^2, the P = rho.g.h = kg/m^3 . m/s^2 . m = kg / (m.s^2) So that all works. Hence _your_ job, is convert - h = height, into meters - rho = density, into kg/m^3 and then apply the formula for P, P = rho.gh

OpenStudy (jamesj):

And 1 Pascal = 1 Pa = 1 kg/(m.s^2). So these are the correct units.

OpenStudy (anonymous):

thank you